OpenStudy (anonymous):

let T:R^2 to R^2 be linear operator given by formula T(x,y)=(2x-y,-8x+4y) then which of the following vectors are in kernal(T) a)(5,10) b)(3,2) c)(1,1)

OpenStudy (nikita2):

a

myininaya (myininaya):

its because T(5,10)=(0,0) right?! :)

OpenStudy (anonymous):

if a vector, say v, is in the kernal(T), then that means T(v) = 0, so lets plug in the vectors and see which one comes out to 0: T(5,10) = (2(5)-10, -8(5)+4(10)) = (10-10, -40+40) = (0,0) wow thats luck lol, its a.

OpenStudy (anonymous):

on a side note, i dont like calling it the Kernal, i like calling it the Null Space. Sounds much more awesome lol

OpenStudy (anonymous):

thank you sir i got the answer